OpenStudy (anonymous):

evaluate the ff: 6 5/2 x 6 3/2

OpenStudy (king):

63 3/4

OpenStudy (anonymous):

\[6\frac{5}{2} \times 6\frac{3}{2} = \frac{17}{2} \times \frac{15}{2} = \frac{255}{4} = 63\frac{3}{4}\]